OpenStudy (anonymous):

When solving a radical equation, Amari and Justin came to two different conclusions. Amari found a solution, while Justin's solution did not work in the equation. Create and justify two situations: one situation where Amari is correct and a separate situation where Justin is correct.

OpenStudy (anonymous):

@ayeitsJenni. @dwatts158

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

first thing is do u know what the radical equations look like

OpenStudy (anonymous):

How do I find that?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

It's extraneous right?

OpenStudy (anonymous):

\[\sqrt{x-2}=5\]

OpenStudy (anonymous):

Square both sides: x - 2 = 25 x = 25 + 2 = 27 Put x = 27 back in the original equation and see if the solution is correct: sqrt(27-2) = sqrt(25) = 5 which is the same as the right hand side and therefore the solution is correct. Next take the radical equation: ...

OpenStudy (mathmath333):

yes this is an example for radical equation

OpenStudy (anonymous):

So I could use this?

OpenStudy (anonymous):

What's one that won't work?

OpenStudy (mathmath333):

what is \(\Large \sqrt{25}\)

OpenStudy (anonymous):

5

OpenStudy (mathmath333):

what is \(\Large -5\times -5=\)

OpenStudy (anonymous):

25?

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

so actually \(\huge \sqrt{25}=\pm5\)

OpenStudy (mathmath333):

did u get that

OpenStudy (anonymous):

Yea, and?

OpenStudy (mathmath333):

so if u consider \(\sqrt{x-2}=5\) for which u r getting the answer as \(27\) \( \sqrt{27-2}=\sqrt{25}\\= -5\) if u consider the sqrare root of 25 as -5 it wull be wrong according to your original equation cuz \(-5\neq 5\)

OpenStudy (anonymous):

OH!!! That makes sense... THANKS a lot!

OpenStudy (mathmath333):

yw